# Improved Streaming Algorithm for Fair $k$-Center Clustering

### v1

@@ 3.2 Procession of Case (2)
To prove $|\Gamma_{l}^{\prime}|\leq k_{l}$ of Cond. (1), we first show that each $i\in\Gamma_{l}^{\prime}$ can not be covered by any point of $S_{3-l}$, i.e., $d(i,j)>r^{*}$ holds for any point $j\in S_{3-l}$. Suppose this is not true, i.e., for a point $i\in\Gamma_{l}^{\prime}$, there exists a point $j$ in $S_{3-l}$ with $d(i,j)\leq r^{*}$. According to the definition of the $2r^{*}$-independent center set, there must exist a point $p\in\Gamma_{3-l}$ with $d(j,p)\leq 2r^{*}$. So we get $d(i,p)\leq d(i,j)+d(j,p)\leq r^{*}+2r^{*}=3r^{*}$, where the second inequality is by the triangle inequality. This contradicts with the definition of $\Gamma_{l}^{\prime}$ that $d\left(i,\Gamma_{3-l}\right)>3r^{*}$ holds for each $i\in\Gamma_{l}^{\prime}$. Then, $|\Gamma_{l}^{\prime}|\leq k_{l}$ must hold, because otherwise there must exist at least two points $p,q\in\Gamma_{l}^{\prime}$ belonging to an identical cluster in the optimal solution. That means $d(p,q)\leq 2r^{*}$, contradicting the fact $d(p,q)>2r^{*}$ as $p,q$ are two points of the $2r^{*}$-independent center set $\Gamma_{l}$.
@@
Next, we show $d(s,C)\leq 5r^{*}$. Firstly, for each point $s\in S_{l}$, following the algorithm, there exists a point $p\in\Gamma_{l}$ such that $d(s,p)\leq 2r^{*}$. Then if point $p\in\Gamma_{l}^{\prime}$, $d(s,p)\leq 2r^{*}$ is true; otherwise, i.e. $p\notin\Gamma_{l}^{\prime}$, then there exists a point $q\in\Gamma_{3-l}$ with $d(p,q)\leq 3r^{*}$, indicating $d(s,q)\leq d(s,p)+d(p,q)\leq 2r^{*}+3r^{*}=5r^{*}$. Secondly, for each point $s\in S_{3-l}$, there exists a point $p\in\Gamma_{3-l}$ such that $d(s,p)\leq 2r^{*}$. Moreover, $\Gamma_{3-l}$ remains unchanged during the algorithm, so $d(s,p)\leq 2r^{*}$ remains true for $p\in\Gamma_{3-l}$. Therefore, for each $s\in S=S_{1}\cup S_{2}$, $d(s,C)\leq 5r^{*}$ holds. ∎
3.3 Procession of Case (3)
@@
We propose a more sophisticated algorithm to ensure the approximation ratio $5$ for Case (3). The key idea of our algorithm is to construct an auxiliary bipartite graph $G(\Gamma_{1}\cup\Gamma_{2})$, such that the center selection problem regarding $\Gamma_{1}\cup\Gamma_{2}$ is transformed into a special constrained vertex cover problem therein. The construction of $G$ simply proceeds as: (1) Set $V(G)=\Gamma_{1}\cup\Gamma_{2}$; (2) For a pair of $p\in\Gamma_{1}$ and $q\in\Gamma_{2}$, add an edge $e(p,q)\in E(G)$ if and only if $d(p,q)\leq 3r^{*}$ holds.
@@
Then, the aim is to find a vertex cover in $G$ with a size bounded by $k$ that also satisfies the fairness constraints $k_{1}$ and $k_{2}$. Notably, although finding a vertex cover is $NP$-hard in general, we manage to devise a polynomial-time exact algorithm for the problem based on certain special properties of the constructed auxiliary graph.
@@
The key idea of our algorithm is to repeatedly eliminate degree 0 and degree 1 points in $G$ when there exist any, by selecting the point covering such points from $V(G)$ as a new center of the center set $C$. While $G$ contains no degree-0 or degree-1 points, our algorithm arbitrarily chooses an edge in $G$, and selects one of its endpoints as a new center by adding it to $C$. The procedure repeats until the center set $C$ covers all points of $G$. The detailed algorithm is illustrated in Alg. 1.
Input: $\Gamma_{1}$ and $\Gamma_{2}$ with $|\Gamma_{1}|>k_{1}$, $|\Gamma_{2}|>k_{2}$
@@
1
// Phase 1: Construction of the auxiliary graph.
2
3Set $C\leftarrow\emptyset$;
4
5Construct the auxiliary graph $G(\Gamma_{1}\cup\Gamma_{2},E)$, where an edge $e(p,q)$ exists in $E$ if and only if $p\in\Gamma_{1}$, $q\in\Gamma_{2}$ and $d(p,q)\leq 3r^{*}$;
6for each point $i\in G$ with degree 0 do
7 if $d(C,i)>2r^{*}$ then
@@ 8     Set $C\leftarrow C\cup\left\{i\right\}$;
9
10Remove each point $i\in G$ with degree 0;
// Phase 2: Computation of $C$.
@@ 22     Set $C\leftarrow C\cup\left\{i\right\}$ and $G\leftarrow G\setminus\{i\}\setminus N_{1}(i)$;
27 if there exists $l$ with $|C\cap S_{l}|+|\Gamma_{l}|\leq k_{l}$ then
28      Set $C\leftarrow C\cup\Gamma_{l}$ and compute $\Gamma^{\prime}_{3-l}=\{p\mid p\in\Gamma_{3-l}\text{ and $d(p,C)>3r^{*}$}\}$
29     Return $C\leftarrow C\cup\Gamma^{\prime}_{3-l}$.
30
31
Algorithm 1 Procession of Case (3)
@@
Each pair of points of $\Gamma_{l}$ belongs to different clusters of the optimum solution when $\lambda=2r^{*}$ before commencing of Alg. 1.
@@
When $\lambda=2r^{*}$, according to Def. 1, the distance between each pair of points in $\Gamma_{l}$ is larger than $2r^{*}$. Suppose that there exists a pair of points $i$, $j$ in $\Gamma_{l}$ such that they are in the same cluster of the optimum solution, that is, there exists a center $c^{*}$ in the center set of the optimum solution such that $d(i,c^{*})\leq r^{*}$ and $d(j,c^{*})\leq r^{*}$ both hold, then $d(i,j)\leq d(i,c^{*})+d(j,c^{*})\leq r^{*}+r^{*}=2r^{*}$, contradicting to $d(i,j)>2r^{*}$. ∎
@@
For the set $C$, $|C\cap S_{l}|\leq k_{l}$ holds when Phase 1 of Alg. 1 completes.
@@
Following Steps 3-6 in Alg. 1, we add $i\in G$ with degree 0 to $C$. That is, each point $i\in\Gamma_{l}$ added in $C$ satisfies $d(i,\Gamma_{3-l})>3r^{*}$. Then according to Cond. (1) in Thm. 1, $d(i,j)>r^{*}$ holds for any point $j\in S_{3-l}$. That is, each point $i\in C\cap\Gamma_{l}$ can not belong to a cluster centered at a point of $S_{3-l}$ in the optimal solution. So similar to the proof of Cond. (1) of Thm. 1, we have $|C\cap S_{l}|\leq k_{l}$. ∎
For briefness, we use $\Gamma^{j}_{l}$ and $C^{j}$ to respectively denote $\Gamma_{l}$ and $C$ in the $j$th iteration of the while-loop of Alg. 1.
@@
In the $j$th iteration of Alg. 1’s while-loop, we have: (1) $|\Gamma^{j}_{l}|\leq k-|C^{j}|$; (2) $|C^{j}|\leq k$; (3) $|C^{j}\cap S_{l}|\leq k_{l}$.
@@
For Cond. (1), we first show it is true when $j=1$. Before the first while loop commences, $|\Gamma^{1}_{l}|\leq k-|C^{1}|$ is satisfied. Suppose otherwise, there would exist a contradiction as analyzed below. We will show the distance between every two points belonging to $C^{1}\cup\Gamma_{l}^{1}$ is larger than $2r^{*}$, so every two points therein can not belong to the same cluster in an optimal solution, and hence there are at least $|\Gamma^{1}_{l}|+|C^{1}|>k$ centers in an optimal solution, arising a contradiction. This fact is deduced following Steps 3-6 in Alg. 1, where points with a zero degree are added to $C^{1}$ for the first iteration. For each two points $i\in C^{1}\cup\Gamma_{l}\cap S_{l}$ and $j\in C^{1}\cup\Gamma_{l}^{1}\cap S_{3-l}$, the absence of edge $e(i,j)$ in $G$ indicates $d(i,j)>3r^{*}$ according to the construction of the auxiliary graph $G$. Moreover, for each two points $i,j\in C^{1}\cup\Gamma_{l}^{1}\cap S_{l}$, $d(i,j)>2r^{*}$ holds because $\Gamma_{l}$ is a $2r^{*}$-independent center set for $S_{l}$($l\in\{1,2\}$), and $i,j$ are points in initial $\Gamma_{l}$. Therefore, $d(i,j)>2r^{*}$ holds for each two points $i,j\in C^{1}\cup\Gamma_{l}$.
We demonstrate that Cond. (1) holds for the $(j+1)$th iteration by induction. Assuming Cond. (1) is valid for the $j$th iteration concerning $\Gamma^{j}$, we either remove two points or remove a point $i$ along with its degree-1 neighbors, each belonging to $\Gamma^{j}_{1}$ and $\Gamma^{j}_{2}$ respectively. Consequently, after this removal, it follows that:
@@
Moving on to Cond. (3), Lem. 3 establishes that in Steps 3-6 of Alg. 1, the inequality $|C\cap\Gamma_{l}|=\gamma_{l}\leq k_{l}$ is valid. That is, Cond. (3) holds for $j=1$. Our objective now is to verify that this inequality persists throughout the $(j+1)$th iteration of Alg. 1. Assume that Cond. (3) holds for $j$th iteration. According to Steps 8-15, Alg. 1 involves adding a single point to $C$ in each iteration. If the algorithm does not terminate in the $(j+1)$th iteration, the inequality $|C^{j}\cap S_{l}|\leq k_{l}$ holds for each $l\in\{1,2\}$ clearly. Otherwise, the algorithm terminates in the $(j+1)$th iteration. Upon the termination of Alg. 1, we encounter two cases: (1) $|C^{j}\cap S_{l}|+|\Gamma_{l}|\leq k_{l}$ and $|C^{j}\cap S_{3-l}|<k_{3-l}$; (2) $|C^{j}\cap S_{3-l}|=k_{3-l}$. Clearly, the inequality $|C^{j}\cap S_{l}|\leq k_{l}$ holds for each $l\in\{1,2\}$ when Case (1) occurs. Regarding the latter Case (2), when a point from $\Gamma_{3-l}$ is added to $C^{j}$ in the $(j+1)$th iteration, we achieve $|C^{j}\cap S_{3-l}|=k_{3-l}$. In accordance with Cond. (1), this leads to $|C^{j}\cap S_{l}|+|\Gamma_{l}|\leq k-|C^{j}\cap S_{3-l}|=k-k_{3-l}=k_{l}$. Therefore, the inequality $|C^{j}\cap S_{l}|\leq k_{l}$ is consistently upheld for each $l\in\{1,2\}$ during the execution of the entire algorithm. This completes the proof. ∎
@@
Alg. 1 terminates in runtime $O(k^{2})$ and outputs a feasible solution $C$ to fair $k$-center, for which $d(s,C)\leq 5r^{*}$ holds for any point $s\in S$.
@@
For the runtime, Step 2 of Alg. 1 takes $O(k^{2})$ to construct $G$ as $G$ contains at most $O(k)$ vertices and $O(k^{2})$ edges. Then, the while-loop (Steps 7-17) iteration for at most $O(k)$ times as it removes at least one points from $G$, and each iteration takes $O(k)$ time. Therefore, the total runtime sums up to $O(k^{2})$.
@@
Therefore, regardless $i\in C$ holds or not, $d(s,C)\leq 5r^{*}$ holds for any $s\in S_{l}$ for any ll.∎
Combining Case (1), (2) and (3), we eventually achieve a complete streaming algorithm for fair $k$-center provided that $r^{*}$ is known. For the space complexity, the algorithm stores $m=2$ independent center sets for each case where each independent center set has $O(k)$ points, so it consumes a space complexity of $O(k)$. We next analyze the update time of the algorithm. In the streaming stage, upon the arrival of each point $i$, the algorithm needs to verify its distance to every point within the current independent center sets, which contains $O(k)$ points. So the update time is $O(k)$. Combining Thm. 2, we have:
@@
Provided $r^{*}$ is known, fair $k$-center admits a streaming algorithm that achieves a ratio 5, consumes $O(k)$ memory and $O(k)$ update time.
Moreover, as $r^{*}$ is actually unknown, we need a $O(\log n)$ multiplicative factor over the memory complexity and overall runtime (Guo et al. 2025). So the algorithm consumes $O(k\log n)$ memory and $O(nk\log n)$ runtime in total.
Notably, the approximation ratio 5 is tight according to the example in the appendix. We can easily extend the algorithms respectively for Cases (2) and Cases (3) and obtain an algorithm for general $m$. Moreover, the streaming algorithm can be tuned to approximate the offline fair $k$-center with a ratio $3$ (shown in the appendix).

### v2

@@ 3.2 Processing of Case (2)
To prove $|\Gamma_{l}^{\prime}|\leq k_{l}$ of Cond. (1), we first show that each $i\in\Gamma_{l}^{\prime}$ cannot be covered by any point of $S_{3-l}$ in the optimal clustering, i.e., $d(i,j)>r^{*}$ holds for any point $j\in S_{3-l}$. Suppose this is not true, i.e., for a point $i\in\Gamma_{l}^{\prime}$, there exists a point $j$ in $S_{3-l}$ with $d(i,j)\leq r^{*}$. According to the definition of the $2r^{*}$-independent center set, there must exist a point $p\in\Gamma_{3-l}$ with $d(j,p)\leq 2r^{*}$. So we get $d(i,p)\leq d(i,j)+d(j,p)\leq r^{*}+2r^{*}=3r^{*}$, where the second inequality is by the triangle inequality. This contradicts the definition of $\Gamma_{l}^{\prime}$ that $d\left(i,\Gamma_{3-l}\right)>3r^{*}$ holds for each $i\in\Gamma_{l}^{\prime}$. Then, $|\Gamma_{l}^{\prime}|\leq k_{l}$ must hold, because otherwise there must exist at least two points $p,q\in\Gamma_{l}^{\prime}$ belonging to an identical cluster in the optimal solution. That means $d(p,q)\leq 2r^{*}$, contradicting the fact $d(p,q)>2r^{*}$ as $p,q$ are two points of the $2r^{*}$-independent center set $\Gamma_{l}$.
@@
Next, we show $d(s,C)\leq 5r^{*}$ for any $s\in S$. Firstly, following the algorithm, there exists a point $p\in\Gamma_{l}$ for any $s\in S_{l}$ such that $d(s,p)\leq 2r^{*}$. Then if point $p\in\Gamma_{l}^{\prime}$, $d(s,p)\leq 2r^{*}$ is true; otherwise, i.e. $p\notin\Gamma_{l}^{\prime}$, then there exists a point $q\in\Gamma_{3-l}$ with $d(p,q)\leq 3r^{*}$, indicating $d(s,q)\leq d(s,p)+d(p,q)\leq 2r^{*}+3r^{*}=5r^{*}$. Secondly, for each point $s\in S_{3-l}$, there exists a point $p\in\Gamma_{3-l}$ such that $d(s,p)\leq 2r^{*}$. Moreover, $\Gamma_{3-l}$ remains unchanged during the algorithm, so $d(s,p)\leq 2r^{*}$ remains true for $p\in\Gamma_{3-l}$. Hence, for each $s\in S=S_{1}\cup S_{2}$, $d(s,C)\leq 5r^{*}$ holds. ∎
3.3 Processing of Case (3)
@@
We propose a more sophisticated algorithm to ensure the approximation ratio of $5$ for Case (3). The key idea of our algorithm is to construct an auxiliary bipartite graph $G(\Gamma_{1}\cup\Gamma_{2},E)$, such that the center selection problem regarding $\Gamma_{1}\cup\Gamma_{2}$ is transformed into a special constrained vertex cover problem in $G$. The construction of $E(G)$ simply proceeds as:
  • For every pair of points $p\in\Gamma_{1}$ and $q\in\Gamma_{2}$, add an
    edge $e(p,q)$ to $E(G)$ if and only if $d(p,q)\leq 3r^{*}$.
@@
Then, the aim is to find a vertex cover in $G$ with a size bounded by $k$ that also satisfies the fairness constraints $k_{1}$ and $k_{2}$. Notably, although finding a vertex cover is $\mathcal{NP}$-hard in general, we manage to devise a polynomial-time exact algorithm for the problem based on certain special properties of the constructed auxiliary graph.
@@
The key idea of our algorithm is to repeatedly eliminate degree 0 and degree 1 points in $G$ when any exist, by selecting the point covering such points from $V(G)$ as a new center in the set $C$. While $G$ contains no degree-0 or degree-1 points, our algorithm arbitrarily chooses an edge in $G$, and selects one of its endpoints as a new center by adding it to $C$. The procedure repeats until the problem reduces to Case (2). The detailed algorithm is illustrated in Alg. 1.
Input: $\Gamma_{1}$ and $\Gamma_{2}$ with $|\Gamma_{1}|>k_{1}$, $|\Gamma_{2}|>k_{2}$.
@@
// Phase 1: Construction of $G$ and $C$.
1 Set $C\leftarrow\emptyset$;
2 Construct the auxiliary graph $G(\Gamma_{1}\cup\Gamma_{2},E)$, where an edge $e(p,q)$ exists in $E$ if and only if $p\in\Gamma_{1}$, $q\in\Gamma_{2}$ and $d(p,q)\leq 3r^{*}$;
3 for each point $i\in G$ with degree 0 do
4 if $d(i,C)>2r^{*}$ then
@@ 5      Set $C\leftarrow C\cup\left\{i\right\}$;
6
7
8Remove each point $i\in G$ with degree 0;
// Phase 2: Final computation of $C$.
@@ 17      Set $C\leftarrow C\cup\left\{i\right\}$ and $G\leftarrow G\setminus\{i\}\setminus N_{1}(i)$;
19
20
21 if there exists $l$ with $|C\cap S_{l}|+|\Gamma_{l}|\leq k_{l}$ then
22
23     Call the algorithm for Case (2) to find the center set $C^{\prime}$ on $\Gamma_{l}\cup\Gamma_{3-l}$ with $k^{\prime}_{l}=k_{l}-|C\cap S_{l}|$.
24     Return $C\cup C^{\prime}$.
@@
Algorithm 1 Processing of Case (3)
@@
Each pair of points from the same $\Gamma_{l}$ belongs to different clusters of the optimum solution when $\lambda=2r^{*}$ before commencing of Alg. 1.
@@
When $\lambda=2r^{*}$, according to Def. 1, the distance between each pair of points in $\Gamma_{l}$ is larger than $2r^{*}$. Suppose there exists a pair of points $i,j$ in $\Gamma_{l}$ such that they are in the same cluster of the optimum solution. That is, there exists a center $c^{*}$ in the center set of the optimum solution such that $d(i,c^{*})\leq r^{*}$ and $d(j,c^{*})\leq r^{*}$ both hold. So $d(i,j)\leq d(i,c^{*})+d(j,c^{*})\leq r^{*}+r^{*}=2r^{*}$, contradicting $d(i,j)>2r^{*}$. ∎
@@
When Phase 1 of Alg. 1 completes, $|C\cap S_{l}|\leq k_{l}$ holds for the current center set $C$.
@@
Following Steps 3-6 in Alg. 1, we add $i\in G$ with degree 0 to $C$. That is, each point $i\in\Gamma_{l}$ added to $C$ satisfies $d(i,\Gamma_{3-l})>3r^{*}$. Then for any point $j\in S_{3-l}$, we have $d(j,\Gamma_{3-l})\leq 2r^{*}$, and hence $d(i,j)>r^{*}$ holds. That is, each point $i\in C\cap\Gamma_{l}$ cannot belong to a cluster centered at a point of $S_{3-l}$ in the optimal solution. On the other hand, each pair of points $i,j\in C$ belongs to different optimal clusters because $d(i,j)>2r^{*}$ (similar to the proof of Lem. 2). Therefore, we have $|C\cap S_{l}|\leq k_{l}$ for every $l$. ∎
For brevity, $\Gamma^{j}_{l}$ and $C^{j}$ respectively denote $\Gamma_{l}$ and $C$ in the $j$th iteration of the while-loop (Phase 2) of Alg. 1.
@@
In the $j$th iteration of the while loop of Alg. 1, we have: (1) $|\Gamma^{j}_{l}|\leq k-|C^{j}|$; (2) $|C^{j}|\leq k$; (3) $|C^{j}\cap S_{l}|\leq k_{l}$.
@@
For Cond. (1), we first show it is true when $j=1$. Before the first while loop commences, $|\Gamma^{1}_{l}|\leq k-|C^{1}|$ is satisfied. Suppose otherwise, there would exist a contradiction as analyzed below. We will show the distance between every two points belonging to $C^{1}\cup\Gamma_{l}^{1}$ is larger than $2r^{*}$, so every two points therein cannot belong to the same cluster in an optimal solution, and hence there are at least $|\Gamma^{1}_{l}|+|C^{1}|>k$ centers in an optimal solution, leading to a contradiction. This can be deduced following Steps 3-6 in Alg. 1, where points with a zero degree are added to $C^{1}$ for the first iteration. For each pair of points $i\in C^{1}$ and $j\in\Gamma_{l}^{1}$, the absence of edge $e(i,j)$ in $G$ indicates $d(i,j)>3r^{*}$ according to the construction of the auxiliary graph $G$. Moreover, for each two points $i,j\in\Gamma_{l}^{1}$, $d(i,j)>2r^{*}$ holds because $\Gamma_{l}$ is a $2r^{*}$-independent center set for $S_{l}$ ($l\in\{1,2\}$), and $i,j$ are points in initial $\Gamma_{l}$. Lastly, for each $i,j\in C^{1}$, $d(i,j)>2r^{*}$ holds because of Step 4 of the algorithm. Therefore, $d(i,j)>2r^{*}$ holds for each two points $i,j\in C^{1}\cup\Gamma_{l}$.
We demonstrate that Cond. (1) holds for the $(j+1)$th iteration by induction. Assume that Cond. (1) is valid for the $j$th iteration concerning $\Gamma^{j}$. Then, we either remove two points or remove a point $i$ along with its degree-1 neighbors. Consequently, after this removal, it follows that:
@@
Moving on to Cond. (3), Lem. 3 establishes that in Steps 3-6 of Alg. 1, the inequality $|C^{1}\cap S_{l}|\leq k_{l}$ is true. That is, Cond. (3) holds for $j=1$. Assume that Cond. (3) holds for $j$th iteration. We need only to verify that this inequality holds in the $(j+1)$th iteration of Alg. 1. According to Steps 8-15, Alg. 1 involves adding a single point to $C$ in each iteration. So if $|C^{j}\cap S_{l}|<k_{l}$ holds for any $l\in\{1,2\}$, then $|C^{j+1}\cap S_{l}|\leq k_{l}$. Otherwise (i.e., $|C^{j}\cap S_{l}|=k_{l}$), we have $|\Gamma^{j}_{3-l}|+|C^{j}|\leq k$ by Cond. (1). This implies
$|\Gamma^{j}_{3-l}|+|C^{j}\cap S_{3-l}|\leq k-|C^{j}\cap S_{l}|=k-k_{l}=k_{3-l}.$
According to the if condition in Step 16 of Alg. 1, when the above inequality holds, Alg. 1 terminates and reduces to Case (2). This completes the proof. ∎
@@
Alg. 1 terminates in time $O(k^{2})$ and outputs a feasible solution $C$ to fair $k$-center, for which $d(s,C)\leq 5r^{*}$ holds for any point $s\in S$.
@@
For the runtime, Step 2 of Alg. 1 takes $O(k^{2})$ time to construct $G$ as $G$ contains at most $O(k)$ vertices and $O(k^{2})$ edges. Then, the while-loop (Steps 7-18) iterates for at most $O(k)$ times as it removes at least one point from $G$ in each iteration, and each iteration takes $O(k)$ time. Therefore, the total runtime sums up to $O(k^{2})$.
@@
Therefore, regardless of whether $i\in C$ holds or not, $d(s,C)\leq 5r^{*}$ holds for any $s\in S_{l}$ for any $l\in\{1,2\}$.∎
Combining Cases (1), (2) and (3) with the streaming stage, we eventually achieve a complete streaming algorithm for fair $k$-center provided that $r^{*}$ is known. For the space complexity, the algorithm stores $m=2$ independent center sets for each case where each independent center set has $O(k)$ points, so it consumes a space complexity of $O(k)$. However, as $r^{*}$ is actually unknown, we need an $O(\log\alpha)$ multiplicative factor over the memory complexity, where $\alpha$ is the aspect ratio that is the ratio between the maximum and minimum distance between the points of $S$. Combining Thm. 2, we have:
@@
When $r^{*}$ is unknown, fair $k$-center admits a streaming algorithm with an approximation ratio $5+\epsilon$ and a memory complexity $O(k\log\alpha)$.
Notably, we can extend the algorithms respectively for Cases (2) and Cases (3) and obtain an algorithm for general $m$ (See the extended algorithm in the full version).